# RED-DiffEq: Regularization by denoising diffusion models for solving inverse PDE problems with application to full waveform inversion

### v1

@@ 4.2 Evaluation metrics
RMSE $=\sqrt{\frac{1}{N}\,\sum_{i=1}^{N}\bigl(\mathbf{x}_{i}-\hat{\mathbf{x}}_{i}\bigr)^{2}},$
@@
$\text{SSIM}(\mathbf{x},\hat{\mathbf{x}})\;=\;\frac{\bigl(2\mu_{x}\mu_{\hat{x}}+C_{1}\bigr)\bigl(2\sigma_{x\hat{x}}+C_{2}\bigr)}{\bigl(\mu_{x}^{2}+\mu_{\hat{x}}^{2}+C_{1}\bigr)\bigl(\sigma_{x}^{2}+\sigma_{\hat{x}}^{2}+C_{2}\bigr)},$
@@ Acknowledgments
$\frac{1}{\mathbf{x}^{2}(\mathbf{r})}\frac{\partial^{2}\mathbf{u}(\mathbf{r},t)}{\partial t^{2}}-\nabla^{2}\mathbf{u}(\mathbf{r},t)=q(\mathbf{r},t),$
@@
$\arg\min_{\mathbf{x}}\left[\frac{1}{2}\|\mathbf{u}_{\text{data}}-f_{\text{PDE}}(\mathbf{x})\|_{2}^{2}+\lambda R(\mathbf{x})\right],$
@@
Appendix S2 RED-DiffEq algorithm
Algorithm S1 RED-DiffEq for full waveform inversion.
1: Initialize velocity model: $\mathbf{x}_{0}$
2:for $k=0,\ldots,K-1$ do
3:  $t\sim\mathcal{U}\{1:T\}$ {Sample timestep uniformly (integer)}
9:end for
10:return $\mathbf{x}_{K}$ {Final velocity model}
In our experiment, we adopt the Adam optimizer for implementing gradient descent. The noise schedule $\{\gamma(t)\}_{t=1}^{T}$ specifies the cumulative signal‐ratio at each diffusion step $t=1,\dots,T$. In diffusion models, the noise schedule dictates how noise is progressively added to the input data over $T$ time steps, transforming it into pure noise in the forward process. This schedule is critical for training the model to reverse the process and recover the original signal. In our implementation, we adopt a sigmoid‑based $\gamma$–schedule [68] (Algorithm S2 and Fig. S1).
@@
$s_{\text{data},t}(\mathbf{z})\coloneq\nabla_{\mathbf{z}}\log p_{\text{data},t}(\mathbf{z}).$
@@
$\mathbb{E}\!\big[\mathbf{x}\,\big|\,\mathbf{y}\big]\;=\;\mathbf{y}+\sigma^{2}\nabla_{\mathbf{y}}\log p(\mathbf{y}).$
@@
$\mathcal{D}_{\theta}(\mathbf{x};t,\boldsymbol{\epsilon})\;=\;\frac{1}{\sqrt{\gamma(t)}}\Big(\mathbf{x}_{t}\;-\;\sqrt{1-\gamma(t)}\,\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{t},t)\Big),\quad\mathbf{x}_{t}=\sqrt{\gamma(t)}\,\mathbf{x}+\sqrt{1-\gamma(t)}\,\boldsymbol{\epsilon}.$
@@
The diffusion model was trained on the OpenFWI dataset [62], encompassing FF-B, FV-B, CV-B, and CF-B velocity model families, using a U-Net architecture with ResBlock and Attention Block components (Fig. 2a). The model architecture incorporated 1,000 diffusion steps with a sigmoidal noise schedule (start = 3-3, end = 3, $\tau=1$) in Alg. S2, 64 base channels with progressive channel multipliers (1, 2, 4, 8), and 4 attention heads. We trained the model for 400,000 iterations using a batch size of 32 and a learning rate of 0.0002.
@@
To ensure fair comparison, hyperparameters were kept consistent across all experiments. The optimization process used a learning rate of 0.03 with a cosine annealing scheduler for 300 iterations. Regularization coefficients were tuned through empirical validation: 0.01 for both Tikhonov and Total Variation methods, and 0.7 for RED-DiffEq. These values were selected to achieve a balance between quantitative accuracy and qualitative fidelity, ensuring the regularization terms effectively constrained the solution space while avoiding common artifacts (such as the staircase effect) that compromise geological interpretation.
S4.1 Tikhonov regularization
@@
S4.2 Total variation regularization
@@
Appendix S5 Statistical test of performance for missing traces
To more rigorously assess the impact of missing traces on each method, we conducted a two-way ANOVA for three evaluation metrics: RMSE, MAE, and SSIM. The $p$-values associated with the number of missing traces (5–20) were 0.456, 0.568, and 0.624 for RMSE, MAE, and SSIM, respectively—well above the 0.05 significance threshold—indicating no statistically significant effect of missing traces alone.

### v2

@@ 4.2 Evaluation metrics
RMSE $=\sqrt{\frac{1}{N}\,\sum_{i=1}^{N}\bigl(\mathbf{x}_{i}-\hat{\mathbf{x}}_{i}\bigr)^{2}},$ (19)
@@
$\text{SSIM}(\mathbf{x},\hat{\mathbf{x}})\;=\;\frac{\bigl(2\mu_{x}\mu_{\hat{x}}+C_{1}\bigr)\bigl(2\sigma_{x\hat{x}}+C_{2}\bigr)}{\bigl(\mu_{x}^{2}+\mu_{\hat{x}}^{2}+C_{1}\bigr)\bigl(\sigma_{x}^{2}+\sigma_{\hat{x}}^{2}+C_{2}\bigr)},$ (20)
@@ Acknowledgments
$\frac{1}{\mathbf{x}^{2}(\mathbf{r})}\frac{\partial^{2}\mathbf{u}(\mathbf{r},t)}{\partial t^{2}}-\nabla^{2}\mathbf{u}(\mathbf{r},t)=q(\mathbf{r},t),$ (S1)
@@
$\arg\min_{\mathbf{x}}\left[\frac{1}{2}\|\mathbf{u}_{\text{data}}-f_{\text{PDE}}(\mathbf{x})\|_{2}^{2}+\lambda R(\mathbf{x})\right],$ (S2)
@@
Appendix S2 Noise schedule algorithm
The noise schedule $\{\gamma(t)\}_{t=1}^{T}$ specifies the cumulative signal‐ratio at each diffusion step $t=1,\dots,T$. In diffusion models, the noise schedule dictates how noise is progressively added to the input data over $T$ time steps, transforming it into pure noise in the forward process. This schedule is critical for training the model to reverse the process and recover the original signal. In our implementation, we adopt a sigmoid‑based $\gamma$–schedule [23] (Algorithm S2 and Fig. S1), as it demonstrates superior generation performance compared to other noise schedules (e.g., linear and cosine).
@@
$s_{\text{data},t}(\mathbf{z})\coloneq\nabla_{\mathbf{z}}\log p_{\text{data},t}(\mathbf{z}).$ (S3)
@@
$\mathbb{E}\!\big[\mathbf{x}\,\big|\,\mathbf{y}\big]\;=\;\mathbf{y}+\sigma^{2}\nabla_{\mathbf{y}}\log p(\mathbf{y}).$ (S5)
@@
$\mathcal{D}_{\theta}(\mathbf{x};t,\boldsymbol{\epsilon})\;=\;\frac{1}{\sqrt{\gamma(t)}}\Big(\mathbf{x}_{t}\;-\;\sqrt{1-\gamma(t)}\,\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{t},t)\Big),\quad\mathbf{x}_{t}=\sqrt{\gamma(t)}\,\mathbf{x}+\sqrt{1-\gamma(t)}\,\boldsymbol{\epsilon}.$ (S7)
@@
The diffusion model was trained on the OpenFWI dataset [12], encompassing FF-B, FV-B, CV-B, and CF-B velocity model families, using a U-Net architecture with ResBlock and Attention Block components (Fig. 2a). The model architecture incorporated 1,000 diffusion steps with a sigmoidal noise schedule (start = 3-3, end = 3, $\tau=1$) in Alg. S2, 64 base channels with progressive channel multipliers (1, 2, 4, 8), and 4 attention heads. We trained the model for 600,000 iterations (parameter update steps over mini-batches) using a batch size of 32 and a learning rate of 0.0002.
The training loss converged after approximately 300,000 iterations and continued to oscillate (Fig. S2, left), which is typical behavior for diffusion model training. We evaluated the Fréchet Inception Distance (FID) score to assess generation quality every 100,000 iterations (Fig. S2, right). While the FID score generally decreases as training progresses, we found that optimal inversion performance does not necessarily align with the lowest FID. Specifically, inversion accuracy is the highest at around 400,000 to 500,000 iterations for both clean seismic data and data contaminated with Gaussian noise ($\sigma=0.5$) (Fig. S3). Therefore, all the diffusion-model-based results reported in this paper are produced using the network trained with 400,000 iterations.
Figure S2: Training dynamics of the diffusion model. (Left) Training loss of the diffusion model. (Right) Fréchet Inception Distance (FID) evaluated during training.
Figure S3: Inversion performance of diffusion models trained with different iterations.
S4.2 Hyper-parameters tuning and design choices
@@
To ensure fair comparison, hyperparameters were kept consistent across all experiments. The optimization process used a learning rate of 0.03 with a cosine annealing scheduler for 300 iterations, we adopt the Adam optimizer for implementing gradient descent on the velocity model, since in RED-DiffEq the optimization objective incorporates a diffusion-based regularization term that introduces stochasticity through random sampling during each iteration, which makes methods like L-BFGS a less optimal choice. Regularization coefficients were tuned through empirical validation: 0.01 for both Tikhonov and Total Variation methods, and 0.75 for RED-DiffEq. These values were selected to achieve a balance between quantitative accuracy and qualitative fidelity, ensuring the regularization terms effectively constrained the solution space while avoiding common artifacts (such as the staircase effect) that compromise geological interpretation.
For the baseline methods DiffusionFWI and DiffusionILVR, we utilized the official implementations from their respective repositories, adapting the source code to support acoustic FWI (replacing the original elastic engine) and to track intermediate inversion results. Hyperparameters were separately fine-tuned on subsets of the OpenFWI and Marmousi datasets. While the authors of DiffusionFWI and DiffusionILVR proposed several stabilization techniques, including gradient smoothing, gradient normalization, and velocity model blurring, our ablation studies indicated that combining gradient smoothing with velocity model blurring yielded the best results in our setting. Consequently, all reported results for DiffusionFWI and DiffusionILVR incorporate these two techniques.
S4.3 Time dependent weight ablation study
We empirically found that replacing the time-dependent weighting $w(t)$ with a fixed parameter $\lambda$ improves performance. Specifically, using a fixed regularization strength yielded an SSIM of 0.7920, MAE of 0.1114, and RMSE of 0.1845. In contrast, the time-weighted approach resulted in an SSIM of 0.7677, MAE of 0.1275, and RMSE of 0.1979. Consequently, the fixed-weight strategy achieves a relative improvement of approximately 3.06% in SSIM, 14.43% in MAE, and 7.27% in RMSE compared to the time-weighted baseline.
S4.4 Forward solver details
The 2D acoustic wave equation is solved using a 4th-order finite difference scheme with a Ricker wavelet source. Sources and receivers are uniformly distributed along the surface. The solver configuration is identical to the one used to generate the OpenFWI dataset (Table S1).
Table S1: Forward solver configuration.
Parameter Symbol OpenFWI Marmousi/Overthrust Unit
Horizontal grid points $n_{x}$ 70 190
Vertical grid points $n_{z}$ 70 70
Grid spacing $\Delta x$ 10.0 10.0 m
Time steps $n_{t}$ 1000 1000
Time step size $\Delta t$ 0.001 0.001 s
Recording time $T$ 1.0 1.0 s
Source frequency (Ricker) $f$ 15.0 15.0 Hz
Number of sources $n_{s}$ 5 5
Number of receivers $n_{g}$ 70 190
Source depth $z_{s}$ 10 10 m
Receiver depth $z_{g}$ 10 10 m
Boundary condition Sponge layer (quadratic damping)
ABC layer thickness $n_{\text{bc}}$ 120 120 grid points
Physical domain size $700\times 700$ $700\times 1900$ m2
Seismic data shape $(5,1000,70)$ $(5,1000,190)$
S4.5 Computational cost for RED-DiffEq
The runtime of RED-DiffEq is summarized in Table S2. RED-DiffEq introduces only a negligible overhead compared to standard FWI without regularization. The additional diffusion step contributes a minor cost, resulting in an overall runtime increase of just 0.8% per iteration. This indicates that incorporating RED-DiffEq has minimal impact on wall-clock time.
In terms of memory usage, RED-DiffEq requires moderately higher GPU memory due to the storage of diffusion model parameters. The peak VRAM usage increases from 5128 MB to 5884 MB, corresponding to a 14.7% overhead. However, this increase remains well within the capacity of modern GPUs and does not represent a practical limitation. Overall, RED-DiffEq achieves its regularization benefits with minimal additional computational and memory costs.
[…]
Process Avg. Time (s)
No Regularization Forward solver 0.7651
Differentiation 2.0306
Total 2.7958
RED-DiffEq Forward solver 0.7643
Diffusion step 0.0162
Differentiation 2.0368
Total 2.8174
Overhead +0.8%
S4.6 Tikhonov regularization
@@
S4.7 Total variation regularization
@@
Appendix S5 Effect of the post-processing refinement step
We evaluate a diffusion-based post-processing refinement step that denoises the reconstructed velocity model using the pretrained diffusion prior, notably without enforcing data fidelity constraints. Starting from the final inverted model, we apply a forward diffusion process up to a chosen timestep $t$, followed by the conventional DDPM reverse process to obtain a refined estimate. While increasing $t$ generally strengthens the prior influence and yields visually cleaner and more geologically plausible structures (Fig. S4, left), the absence of a data-consistency term implies that an overly large $t$ can bias the solution away from the ground truth. This trade-off is reflected in the modest degradation of quantitative metrics (RMSE, MAE, SSIM) observed at higher $t$ values (Fig. S4, right). Because this refinement step does not enforce data fidelity, it inherently drives the velocity model toward high-probability structures under the learned prior. This strong prior influence can introduce biased features, thereby increasing quantitative errors.
Figure S4: The effect of the post-processing step across different diffusion timesteps ($t$). “w/o” denotes the result from RED-DiffEq without any post-processing. The variable $t$ indicates the starting diffusion timestep selected for denoising.